How to convert 30.10 in/hg ---> atm. Step by step instructions.
Answer #1

Here, I am considering 'in' as inches.

1 inches = 25.4 mm (milimeters)

So, 30.10 inches = 30*25.4 = 764.54 mm

We know that,

760 mmHg = 1 atm

So, 764.54 mmHg = 764.54/760 = 1.006 atm .... Answer